2. A helium-neon laser (2 633 nm) illuminates a single slit, and a diffraction pattern is observed on a wall 1.5 m behind the slit. The distance between the first and second dark fringes is 4.75 mrm What is the width of the slit? c. 0.35 mm e. 0.004 mm 3. Of the following situations, which one is impossible? a. A body having westward velocity and westward acceleration b. A body having westward velocity and eastward acceleration c. A body having zero velocity and non-zero acceleration d. A body having constant acceleration and changing velocity e. A body having constant velocity and changing acceleration
Answer #1

2] By single slit diffraction formula, lambda/d = y/D

633e-9/d = 4.75e-3/1.5

d = 633e-9/[4.75e-3/1.5] = 0.0002 m = 0.2 mm

option d is correct.